# PRAGMA: A Profiling-Reasoned Multi-Agent Framework for Automatic Kernel Optimization

### v1

@@ Abstract
Designing high-performance kernels requires expert-level tuning and a deep understanding of hardware characteristics. Recent advances in large language models (LLMs) have enabled automated kernel generation, yet most existing systems rely solely on correctness or execution time feedback, lacking the ability to reason about low-level performance bottlenecks. In this paper, we introduce PRAGMA, a profile-guided AI kernel generation framework that integrates execution feedback and fine-grained hardware profiling into the reasoning loop. PRAGMA enables LLMs to identify performance bottlenecks, preserve historical best versions, and iteratively refine code quality. We evaluate PRAGMA on KernelBench, covering GPU and CPU backends. Results show that PRAGMA consistently outperforms baseline AIKG without profiling enabled and achieves 2.81 and 2.30 averaged speedups against Torch on CPU and GPU platforms, respectively.
@@ IV-A Evaluation Setup
(2) The AIKG system (no performance feedback version of PRAGMA).
@@ IV-B Performance: Speedup Distribution
Figure 3 presents the speedup distribution of PRAGMA and AIKG over Torch baseline across the KernelBench dataset. The benchmark kernels are grouped into six categories: matrix multiplication (matmul), activation, normalization, pooling/reduction, convolution (conv), and others (including loss and cumulative-sum kernels).
On the CPU, PRAGMA achieves an average of 2.81 speedup over Torch and 10.95 improvement over AIKG. On the GPU, PRAGMA delivers a 2.30 speedup over Torch and a 1.90 improvement compared with AIKG. Meanwhile, it can be observed that both PRAGMA and AIKG generate GPU kernels that outperform Torch on average. The integration of detailed profiling information allows PRAGMA to achieve a more significant improvement. On the CPU, this effect becomes even more pronounced, as profiling data provides richer insights into memory access patterns and cache utilization which have a greater impact on CPU workloads. These results confirm the effectiveness of our profile-guided feedback loop and iterative optimization mechanism, which enable the system to progressively refine performance.
@@
Although PRAGMA consistently outperforms Torch and AIKG across most categories, we observe limited gains in convolution kernels. Convolution operators exhibit substantially higher implementation complexity, involving deeper loop nests and higher-dimensional tensor accesses. This complexity increases the risk of suboptimal tiling, layout transformation, and prefetching decisions during code generation. As a result, achieving correctness and performance simultaneously often requires multiple iterations, leading to less pronounced acceleration in this category.
@@ IV-C Comparison with Caesar
Table I summarizes the overall results. Caesar achieves a 49.0% Success rate and its generated kernels frequently fail to outperform the original Torch implementations. AIKG improves upon this by introducing multi-agent architecture, achieving 92.0% Success rate, 1.28 average speedup, and 32.2% Fast1. Nevertheless, without profiling-guided insights, AIKG’s optimization remains largely heuristic and insensitive to architectural bottlenecks.
In contrast, PRAGMA introduces a performance-guided multi-agent framework that integrates profiling data from Nsight Compute. This enables fine-grained bottleneck classification and targeted optimization reasoning. As shown in Table I, PRAGMA achieves 2.30 average speedup with 59.8% Fast1, consistently outperforming both Caesar and AIKG.
@@ IV-D1 Max reduction over a dimension
PRAGMA achieves steady improvement from 31.41% to 120%. In contrast, AIKG peaks at 85.97% in the third iteration but later drops to 58.06%. The key difference lies in the feedback mechanism. AIKG only returns execution time after each iteration, providing coarse-grained guidance. PRAGMA, in contrast, integrates fine-grained performance metrics such as memory throughput, arithmetic intensity, and occupancy. These metrics enable precise identification of bottlenecks and direct adjustment of reduction strategies. Crucially, the retention of historical best code and its profiling record allows the PRAGMA to judge whether each modification yields positive or negative performance impact, enabling stable, performance-aware evolution rather than blind trial-and-error.

### v2

@@ Abstract
Designing high-performance kernels requires expert-level tuning and a deep understanding of hardware characteristics. Recent advances in large language models (LLMs) have enabled automated kernel generation, yet most existing systems rely solely on correctness or execution time feedback, lacking the ability to reason about low-level performance bottlenecks. In this paper, we introduce PRAGMA, a profile-guided AI kernel generation framework that integrates execution feedback and fine-grained hardware profiling into the reasoning loop. PRAGMA enables LLMs to identify performance bottlenecks, preserve historical best versions, and iteratively refine code quality. We evaluate PRAGMA on KernelBench, covering GPU and CPU backends. Results show that PRAGMA consistently outperforms baseline N-PRAGMA without profiling enabled and achieves 2.81 and 2.30 averaged speedups against Torch on CPU and GPU platforms, respectively.
@@ IV-A Evaluation Setup
(2) N-PRAGMA i.e., PRAGMA w/o performance feedback.
@@ IV-B Performance: Speedup Distribution
Figure 3 presents the speedup distribution of PRAGMA and N-PRAGMA over Torch baseline across the KernelBench dataset. The benchmark kernels are grouped into six categories: matrix multiplication (matmul), activation, normalization, pooling/reduction, convolution (conv), and others (including loss and cumulative-sum kernels).
On the CPU, PRAGMA achieves an average of 2.81 speedup over Torch and 10.95 improvement over N-PRAGMA. On the GPU, PRAGMA delivers a 2.30 speedup over Torch and a 1.90 improvement compared with N-PRAGMA. Meanwhile, it can be observed that both PRAGMA and N-PRAGMA generate GPU kernels that outperform Torch on average. The integration of detailed profiling information allows PRAGMA to achieve a more significant improvement. On the CPU, this effect becomes even more pronounced, as profiling data provides richer insights into memory access patterns and cache utilization which have a greater impact on CPU workloads. These results confirm the effectiveness of our profile-guided feedback loop and iterative optimization mechanism, which enable the system to progressively refine performance.
@@
Although PRAGMA consistently outperforms Torch and N-PRAGMA across most categories, we observe limited gains in convolution kernels. Convolution operators exhibit substantially higher implementation complexity, involving deeper loop nests and higher-dimensional tensor accesses. This complexity increases the risk of suboptimal tiling, layout transformation, and prefetching decisions during code generation. As a result, achieving correctness and performance simultaneously often requires multiple iterations, leading to less pronounced acceleration in this category.
@@ IV-C Comparison with Caesar
Table I summarizes the overall results. Caesar achieves a 49.0% Success rate and its generated kernels frequently fail to outperform the original Torch implementations. N-PRAGMA improves upon this by introducing multi-agent architecture, achieving 92.0% Success rate, 1.28 average speedup, and 32.2% Fast1. Nevertheless, without profiling-guided insights, N-PRAGMA’s optimization remains largely heuristic and insensitive to architectural bottlenecks.
In contrast, PRAGMA introduces a performance-guided multi-agent framework that integrates profiling data from Nsight Compute. This enables fine-grained bottleneck classification and targeted optimization reasoning. As shown in Table I, PRAGMA achieves 2.30 average speedup with 59.8% Fast1, consistently outperforming both Caesar and N-PRAGMA.
@@ IV-D1 Max reduction over a dimension
PRAGMA achieves steady improvement from 31.41% to 120%. In contrast, N-PRAGMA peaks at 85.97% in the third iteration but later drops to 58.06%. The key difference lies in the feedback mechanism. N-PRAGMA only returns execution time after each iteration, providing coarse-grained guidance. PRAGMA, in contrast, integrates fine-grained performance metrics such as memory throughput, arithmetic intensity, and occupancy. These metrics enable precise identification of bottlenecks and direct adjustment of reduction strategies. Crucially, the retention of historical best code and its profiling record allows the PRAGMA to judge whether each modification yields positive or negative performance impact, enabling stable, performance-aware evolution rather than blind trial-and-error.